OpenStudy (anonymous):

Sq root 75/2 + 3 sq root 2/4 = ? Plz show work

OpenStudy (imstuck):

Does that look like this:\[\frac{ \sqrt{75} }{ 2 }+\frac{ 3\sqrt{2} }{ 4 }\]

OpenStudy (anonymous):

make the denominator the same to where both denominators equal to 4

OpenStudy (anonymous):

@Mmarlowe

OpenStudy (anonymous):

are you still here? @Mmarlowe